OpenStudy (anonymous):

Use factoring to solve this equation. x^2+2x-13=2

OpenStudy (scorcher219396):

First, move everything to one side so the equation equals 0 \[x^2+2x-15=0\] Now you can factor You know it's going to look something like this (x - ? )(x + ?) The two x's have to be in front because they have to multiply to x^2 And one sign will be positive, while the other is negative, because when you multiply the two last terms they have to be negative Now to fill in the ?s, they will have to be two numbers which multiply to -15 and add to 2 Look at a chart of the factors of 15 1 and 15, doesn't work 3 and 5 is next, which will work if we think about the signs One has to be negative, and the two added have to equal positive 2 So the three is negative. And you have (x-3)(x+5)=0 Then, set each of those chunks in parentheses equal to 0, so you have x-3=0 which means x=3, and x+5=0 which means x=-5 Your solutions are x=3,-5

OpenStudy (campbell_st):

rewrite the equation as \[x^2 + 2x - 15 = 0\] find the factors of -15 that add to 2... the larger factor is positive then its (x +factor1)(x + factor2) =0 then solve x + factor1 = 0 and x + factor2 = 0 for the solutions hope it helps

OpenStudy (anonymous):

thx guys